Grade 11Physical Chemistry

The wave length of light having wave number 4000 cm-1 is

Dimple Tunuguntla
7 Years agoGrade 11

1 Answer

Arun
7 Years ago
Dear student
 
wavelength = 1/ wave number
 
hence wave length = 1/4000 = 2.5 * 10-4 cm
 
Regards
Arun (askIITians forum expert)